Maria is three times as old as her brother. In four years, Maria will be twice as old as he will be then. How old is Maria's brother now?
Answer by Maths68(1474)

Let
Present age of Maria = m
Present age of Brother = b
Given
Maria is three times as old as her brother
m=3b ..........(1)
In four years means we have to add 4 in their present ages
m+4=b+4
then Maria will be twice as old
m+4=2(b+4)
m+4=2b+8
m=2b+8-4
m=2b+4 .............(2)
Plug in the value of m from (1) to (2)
3b=2b+4
3b-2b=4
b=4
Brother is 4 years old